Unconstraining Using Double Exp Smth (Holts Model) Cumulative Per iod Sales Level Trend Forecast -0.3 8,000+0.7 (20000) 0.216400-0)+0.8 20000 8,000 0 20,000 20,000 2 1,000 | 16,400 | 19,280 | 35,680 K 44,000 78,000 88,000 106,000 29,000 167,000 179,000 92,000 92,000 92,000 = I 6400 + I 9280 4 6 7

What is the forecasted cumulative sales over all 12 Periods?
Answer #1
Level trend forecast
1 8000 0 20000 20000
2 21000 0.3*8000+0.7*20000          =16400 0.2*(16400-0)+0.8*20000                   =19280 (16400+19280) =35680
3 44000 0.3*21000+0.7*19280        =19796 0.2*(19796-16400)+0.8*19280            =16103.2 (19796+16103.2)      =35899.20
4 78000 0.3*44000+0.7*16103.2     =24472.24 0.2*(24472.24-19796)+0.8*16103.2     =13817.81 (24472.24+13817.81)=38290.05
5 88000 0.3*78000+0.7*13817.81   =33072.47 0.2*(33072.47-24472.24)+0.8*13817.81=12774.29 (33072.47+12774.29)=45846.76
6 106000 0.3*88000+0.7*12774.29   =35342 0.2*(35342-33072.47)+0.8*12774.29   =10673.34 (35342+10673.34)    =46015.24
7 129000   0.3*106000+0.7*10673.34 =39271.34 0.2*(39271.34-35342)+0.8*10673.34   =9324.54 (39271.34+9324.54) =48595.88
8 167000 0.3*129000+0.7*9324.54 =45227.18 0.2*(45227.18-39271.34)+0.8*9324.54 =8650.80 (45227.18+8650.80) =53877.98
9 179000 0.3*167000+0.7*8650.80 =56155.56 0.2*(56155.56-45227.18)+0.8*8650.80 =9106.32 (56155.56+9106.32) =65261.88
10 192000 0.3*179000+0.7*9106.32 =60074.42 0.2*(60074.42-56155.56)+0.8*9106.32 =8068.83 (60074.42+8068.83) =68143.25
11 192000 0.3*192000+0.7*8068.83 =63248.18 0.2*(63248.18-60074.42)+0.8*8068.83 =7089.81 (63248.18+7089.81) =70337.99
12 192000 0.3*192000+0.7*7089.81 =62562.87 0.2*(62562.87-63248.18)+0.8*7089.81 =5534.79 (62562.87+5534.79) =68097.66